OpenStudy (anonymous):

The resistance (R) in a wire increases when the length of the wire (L) increases. Write the correct equation for this scenario and solve for when the length is 1. Length 10 resistance 2 15 3 .

OpenStudy (anonymous):

When length is 10, resistance is 2 when length is 15, resistance is 3 according to question \[length \space\ \alpha \space\ resistance\] (length is directly proportional to resistance) \[l \space= k*r\] \[10 \space= k*2\] thus \[k \space= 5\] thus \[l \space= 5*r\] so for length = 1 \[1 \space= 5*r\] \[1/5 \space= r\]